explain the total income of x and y is Rs.9000 x spends 60% of his income and y spends 80% of his income.If their savings are equal ,then the income of x is??
Answer by Theo(13342)

x + y = 9000

x represents the amount of money that person A makes, i.e. his income.
y represents the amount of money that person B makes, i.e. his income.

person A spends 60% of his income.
therefore, what he has left is 40% of his income.
this represent his savings.
this can be shown as .4x

person B spends 80% of his income.
therefore, what he has left is 20% of his income.
this represents his savings.
this can be shown as .2y

you are given that their savings are equal.

.4x = .2y

you have 2 equations that need to be solved simultaneously.

they are:

x + y = 9000
.4x = .2y

this can be solved by substitution fairly easily.

solve for y in the first equation to get:

y = 9000 - x

replace y with 9000 - x in the second equation to get:

.4x = .2 * (9000 - x)

simplify to get:

.4x = 1800 - .2x

add .2x to both sides of this equation to get:

.6x = 1800

divide both sides of this equation by .6 to get:

x = 1800 /.6 = 3000

since x + y = 9000, then y = 6000

you have:

income of person A is 3000
income of person B is 6000

person A spends 60% of 3000 and is therefore left with 40% of 3000 which is equal to 1200.

person B spends 80% of 6000 and is therefore left with 20% of 6000 which is equal to 1200.

their savings are equal.

you are asked for the income of x.

x represents the income for person A which is equal to 3000.